OpenStudy (anonymous):

How would I solve this problem? Explain the appropriate translations of your graphs. 1. f(x)= 2(x + 10)^2 - 2 to f(x)=2(x+8)^2-7

OpenStudy (jdoe0001):

f(x)= 2(x + 10)^2 - 2 ^ ^ horizontal shift vertical shift to f(x)= 2(x+8)^2 -7 ^ ^ horizontal shift vertical shift

OpenStudy (jdoe0001):

y = A ( Bx + C ) + D C = horizontal shift, C > 0, to the left, C < 0, to the right D = vertical shift, D > 0, up, D < 0 down